5 cm. diameter hole is drilled in a 45 cm. by 24 cm. flat rectangular piece of material.
1. Find area of the rectangular piece after the circular hole is removed.
2. What percent of the total area is removed from the rectangular piece?

Answer by Cromlix(4381)

Area of rectangular piece of material
= 45 * 24 = 1080cm^2
Area of hole = Pi*r^2
= 19.64cm^2
Area left after circle removed:-
1080 - 19.64 = 1060.36cm^2
Percentage of total area = 19.64/1080 * 100
= 1.8% is removed.
Hope this helps.
:-)